OpenStudy (anonymous):

If f(x) = log3 (x + 1), what is f−1(2)? 1 8 10 27

hartnn (hartnn):

try to find x from log3 (x+1) = 2 that would be your answer.

hartnn (hartnn):

use the definition off log then \(\Large \log_ab=c \implies b=a^c\)

hartnn (hartnn):

so, \(\Large (x+1)=3^2\) x= ... ?

OpenStudy (anonymous):

Thanks hart!

hartnn (hartnn):

welcome blake ^_^ oh, and \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)